You invested a total of 40,000 at 3% and 4 1/2% simple interest. During one year, the two accounts earned $1575. How much did you invest in each account
Answer by macston(5194)

.
T=amount at 3%; F=amount at 4.5%
.
T+F=$40000
T=40000-F Use to substitute for T.
.
0.03T+0.045F=$1575 Substitute for T.
0.03($40000-F)+0.045F=$1575
$1200-0.03F+0.045F=$1575
0.015F=$375
F=$25000 ANSWER 1: $25000 was invested at 4.5%.
T=$40000-F=$40000-$25000=%15000 ANSWER 2: $15000 was invested at 3%.
.
CHECK:
.
0.03T+0.045F=$1575
0.03($15000)+0.045($25000)=$1575
$450+$1125=$1575
$1575=$1575